# The Relational Origins of Rules in Online Communities

### v1

@@ 3.1. Subreddits
Subreddits are user created forum spaces on the networked community platform Reddit.com. A Reddit user creates a subreddit by giving it a name and a topic. Other users may contribute posts (threads to start off a discussion) or comments (to discuss those posts). Subreddits have been created to support a wide variety of topics, from posting pictures of cute animals to discussing the implications of AI. Reddit users can search for subreddit communities and join by subscribing to a feed that shows the user posts from subreddits they follow.
Governance of subreddits is primarily enacted by the creator of the subreddit and other volunteer content moderators they appoint (Seering et al., 2019). Reddit provides community leaders with a suite of tools, including automated rule enforcement with the AutoModerator bot (Jhaver et al., 2019b), membership management, and a member communication system called Mod Mail. Subreddit leaders often prominently display their rules on the side of the subreddits’ main page and use built-in wiki features to elaborate on community guidelines and FAQ’s. Subreddit rules usually have to do with community topic or behavioral expectations (Fiesler et al., 2018). Studies in social computing have mapped out the kinds of rules that are adopted across subreddits (Chandrasekharan et al., 2018; Fiesler et al., 2018) and quantitative factors for rule change (Reddy and Chandrasekharan, 2023).
@@ 3.2. Fandom wikis
Wikis are collaboration-based websites where users peer produce public information repositories, like an encyclopedia, by directly creating, editing, and organizing content on the website’s pages. Fandom is a platform that affords the creation and management of wikis, often made up of members of a fandom who share a strong interest in documenting information about a specific work of fiction or cultural artifact (e.g., TV show, book series, or video game). Interactions primarily occur on the “Talk Pages” of articles in a Fandom wiki, where members discuss article content. Some Fandom wikis choose to activate a feature in the platform that provides the community with a simple discussion forum.
Governance in Fandom wikis is more structured and often results in hierarchical roles. As the goal of wikis is to produce relevant, high-quality articles documenting their fandoms, leaders frequently develop rules regulating article content. Users who create a Fandom wiki become its de facto leader but frequently appoint other community members into leadership positions. Social computing scholars have documented how wikis engage in ongoing deliberation about how to change their communities’ rules (Keegan and Fiesler, 2017), and shown that as they grow and mature they become more bureaucratic (by adding more rules) (Butler et al., 2008) and oligarchic and closed to newcomers (Shaw and Hill, 2014; Halfaker et al., 2013; TeBlunthuis et al., 2018).
@@ 3.3. Fediverse servers
The Fediverse is an open, decentralized network of servers (also called “instances”) that run the ActivityPub protocol. Users join the Fediverse by registering an account on a specific server of their choice. The server enables its members to share and receive content from other servers, allowing users to interact with fellow server members and the broader network through comments, reposts (“boosts”), likes, and more. We focused on platforms using the Mastodon and Pleroma software, which both offer a microblogging experience similar to systems like Twitter. Servers we observed were generally organized around a shared interest or identity, but members were often not required to post content related to that theme.
Envisioning decentralized social media ecosystems without centralized owners, Fediverse servers using Mastodon and Pleroma are governed exclusively by the servers’ creators, who are referred to as “admins.” Admins control membership through open, closed, or application-based user registration, moderate content through admin panels and back-end access to data, and can block other servers to prevent direct interactions between users of their server and another (known as “defederation”). Rules on Fediverse servers are usually about speech, focused on civility and legality. Although scholarly attention to Fediverse is relatively nascent, researchers have examined migrations to decentralized platforms (Dunbar-Hester, 2024; Cava et al., 2023; He et al., 2023; Nunes, 2023; Jeong et al., 2024), how platforms conceptualize decentralization (Zulli et al., 2020; Gehl and Zulli, 2023; Raman et al., 2019), community governance challenges (Anaobi et al., 2023; La Cava et al., 2024; Bono et al., 2024; Hwang et al., 2025; Zhang et al., 2024), coordination against hate speech (Caelin, 2022), rules on Mastodon platforms (Nicholson et al., 2023), and the effects of defederation on activity (Colglazier et al., 2024).
@@ 4. Methods
This work was a collaboration involving professors and graduate students from [Anonymized for Review]. In the rest of this section, we outline the steps taken by our team to collect and analyze data.
@@ 4.1. Community Sampling and Participant Recruitment
To investigate the origins of rules, we sought a sample of communities that varied in terms of goals and setting. We focused on three different settings introduced in the previous section—subreddits, Fandom wikis, and Fediverse servers. Between 2020 and 2021, we began by building datasets of community rules from each of these types of sociotechnical systems. Because we were initially motivated by observations of rule similarity across communities (Hwang and Shaw, 2022; Fiesler et al., 2018; Reddy and Chandrasekharan, 2023; Nicholson et al., 2023), we computationally clustered communities within each system by their rule language. We then screened clusters to identify a diverse, theoretically relevant sample of active communities to target for interviews. A more detailed description of this process can be found in Appendix Community Sampling Methods.
@@ 4.2. Interviews
We conducted semi-structured interviews via video calls between June and September 2022. All participants were either volunteer moderators or administrators who had familiarity with their community’s rule history and were above the age of 18. Interviews ranged from 30 to 120 minutes in length, averaging 60 minutes. A combination of trained undergraduate research assistants and the paper authors conducted the interviews following an interview protocol that asked participants to describe or tell us stories about their interactions with rules in their communities (see Appendix C). We asked questions about the origins of their rules, how they were enforced, when they changed and why, and about related challenges in governing online communities.
@@ 4.3. Analytical Process
Our analytical process was based on Charmaz’s approach to grounded theory (Charmaz, 2015), thematic analysis (Braun and Clarke, 2006), and collaborative affinity mapping. Small teams of trained research assistants led by four authors conducted line-by-line coding of interview transcriptions using the Taguette software, inductively coding text to break down participant’s experiences into actions and feelings through gerunds (e.g., “relaxing the rules” or “feeling a sense of belonging”). Together, we developed over 4,500 open initial codes. Discussions across the authorship team led to a rough list of 20 themes these codes covered. Guided by this list, three authors then conducted a round of focused coding, a step which merges similar initial codes together and elevates them with broader conceptual labels. This led to 337 unique focused codes.
Three authors then collaboratively engaged in an in-depth affinity mapping exercise with the 337 focused codes to visualize patterns, themes, and processes via a digital whiteboard. Rounds of memo-writing and weekly meetings for discussion and iteration complemented this process. The three authors reached agreement on the most significant contributions to bring forward in this paper: rule adoption and rule change. Finally, the first author engaged in a round of theoretical sampling by using the list of 337 focused codes as a codebook to apply to all 40 interview transcripts. We did this to identify supporting evidence across all three empirical settings and to analyze differences when patterns varied between them. The first author distilled a core set of processes and beliefs around rule adoption and change from this analysis, which was iterated upon with feedback and input from the other authors.
@@ 5.1. Searching for Rule Templates
Few leaders we spoke to imagined a rule set during community creation. Instead, most leaders described the action of searching for rule templates in other online communities. For example, FA14, the leader of a Fandom wiki related to Lego, the plastic construction toy, told us that they “are always inspired by other communities” and “always take a template of rules” from other communities. Some leaders described feelings of not being “actually experienced enough and knowledgeable enough” (FV12) to write rules themselves. FV10, a Fediverse leader for a regional community in the U.S., told us how “a big chunk” of their rules were made at community creation, saying that they “looked to other examples of rules” across Fediverse communities. They referenced their legal background to explain choosing templates over drafting rules themselves:
@@ 5.1.1. Regulating Content and Behavior
In the process for searching for rule templates, leaders primarily cited two motivations: regulating content and regulating behavior. For subreddit leaders like R1, rules regulating content were important for facilitating meaningful discussions. R1 explained why their Korean drama fan community’s rules about mandating substantial discussion posts were important:
@@
Leaders on subreddits often implemented rules on post formatting and content. For example, some communities had strict rules on advertising, such as R4’s cybersecurity discussion community. Others had regulations on content they called “misinformation,” such as R5’s regional community, which had rules adopted to prevent the spread of conspiracies related to the COVID-19 pandemic. We found that Fediverse servers rarely required specific kinds of content in their rules. However, they did restrict kinds of content that are inappropriate for professional work settings (NSFW content), hateful content, and advertising.
Fandom leaders frequently saw their rules as setting the standards for contributions to wikis. For FA6, a leader of a Fandom wiki for a popular multiplayer video game, rules governed information presentation and standards for accuracy. This was typical of rules across the Fandom wikis, almost all of which have content requirements and restrictions on formatting, article relevance, and accuracy.
Behavioral regulation was another important justification for rule adoption across all three settings. For subreddit leaders, this was expressed by a desire to ensure discussion spaces remained civil and safe. R9, a leader of a subreddit that organizes trades of living plants, summarized the sentiment that we heard from other subreddit leaders: “I think the biggest purpose of the rules is to keep everybody safe—to make sure nobody’s getting scammed…to make it so the community doesn’t become a toxic place.”
Although Fandom wikis place less emphasis on facilitating discussion, leaders still adopted rules to promote civility in spaces such as article “Talk Pages” and discussion forums, and in affiliated Discord servers that some Fandom communities used for synchronous chat. FA11 summarized the justification for behavioral rules we heard from other Fandom wiki leaders: “They’re there for safety, and they’re also there to make people feel like they’re included and they belong.”
@@ 5.2. Copying Other Communities’ Rules
Leaders gave various accounts of how they adopted rules for their communities after searching and identifying potential templates. They often mentioned two distinct next steps: copying other communities’ rules and customizing the adopted rule set5.3). One of the original questions motivating our study was “Why do so many online communities have similar rules?” In the previous section, we showed how leaders would seek out other communities’ rule sets to use as starting templates. For many, the next step was to copy. This pattern was most evident in our interviews with members of the Fediverse. FV10 explained this step as “copying the good stuff” after they “look[ed] at other examples of how people have done it.” Leaders would copy and paste rules they felt were a good match for their community and then later “sort of build on it” (FV10) with further customization to their communities’ unique needs. That said, many leaders from the Fediverse would copy the entire rule set of another community and never customize or change it. In explaining these actions, these leaders emphasized a relational function of rules with the goal of establishing external legitimacy among other communities in their decentralized networks.
@@ 5.2.1. Establishing External Legitimacy
A key motivation behind the process of copying other communities’ rules was the desire to project legitimate governance (i.e., the presence of responsible content moderation) and communicate alignment with established norms across an ecosystem of online communities. Fediverse leaders have the power to block (“defederate”) communities that threaten their well-being and goals. These same leaders felt their choices were being observed and judged by other admins of other Fediverse servers that might need to consider whether to defederate at some point. FV10 said this adds “an important function of rules” because it “sends signals to other admins that they know what we are moderating.” This external scrutiny from other admins motivated Fediverse leaders to adopt rules because these rules would signal accountable governance and content moderation. In this way, rules help establish external legitimacy for communities in decentralized networks.
Many Fediverse leaders pointed to mastodon.social’s rule set as a benchmark of legitimate community governance. mastodon.social was one of the first communities on the Fediverse and was created by Eugen Rochko. Rochko is the developer of Mastodon, one of the most widely used platform for hosting Fediverse communities. Fediverse leaders often looked to communities like mastodon.social because it’s “the largest instance” and because it’s “hosted by Eugen” (FV5). FV5 further explained, “You know, if you’re gonna copy rules about a new community, you might as well go right from the source. So I copied most or all of the rules.” The “source” in this case was mastodon.social. FV2 came to a similar conclusion:
@@
The observation that so many other communities on the Fediverse also referenced mastodon.social’s rule set was an important reason that FV2 adopted that particular rule set. From what we can infer from FV2, the widespread adoption of mastodon.social’s rules created a kind of self-reinforcing dynamic—communities that mirror those rule sets signal their alignment with the standards and norms of a reputable, established community.
@@
Subreddit and Fandom wiki leaders did not mention copying other communities rules to signal legitimacy in the way that Fediverse leaders did. That said, some Fandom Wiki leaders like FA3 sought out rules from “other big wikis” to see how others addressed similar issues and used their rules as starting points. For some specific rules related to behavior, subreddit leaders like R4 and R5 mentioned copying guidelines from the “Reddiquette,” which is an informal set of behavioral guidelines written by other Reddit leaders similar to the FFMC.
@@ 5.3. Customizing the Adopted Rule Set
Customizing the adopted rule set was a natural next step for many leaders. After all, the communities from which they found their initial rule sets often differ in their specific goals. Several interviewees from Fandom and Reddit offered insight into this process. FA1, a leader of a Fandom wiki documenting aesthetic fashion and cultural trends, explained that “it would be nice” if every other Fandom wiki had the same rules, but that rules have to be modified to each community’s unique needs: “every community is different.” Leaders often searched for specific kinds of rules from different communities to solve their problems. FA4, the leader of a Fandom wiki cataloging the lore of a popular comic book franchise, described this process in the context of creating a rule on image usage by customizing the image policy from another wiki:
@@
We note that changes were not driven purely based on differences in community topic. Leaders also described trying to improve the language or normative effect of a rule. For example, R4 based their rules on other communities they had led and made them “better to a certain degree.” Sometimes leaders described these edits to other rules as minimal and changing them “a little bit.” (FV12) Other times, the process of rewriting the rules could be “agonizing” (FV11) as leaders struggled to make their rules sound “well written.”

### v2

@@ 3.1. Subreddits
Subreddits are user-created forum spaces on the networked community platform Reddit.com. A Reddit user creates a subreddit by giving it a name and a topic. Other users may contribute posts (threads to start off a discussion) or comments (to discuss those posts). Subreddits have been created to support a wide variety of topics, from posting pictures of cute animals to discussing the implications of AI. Reddit users can search for subreddit communities and join by subscribing to a feed that shows the user posts from subreddits they follow.
Governance of subreddits is primarily enacted by the creator of the subreddit and other volunteer content moderators they appoint (Seering et al., 2019). Reddit provides community leaders with a suite of tools, including automated rule enforcement with the AutoModerator bot (Jhaver et al., 2019b), membership management, and a member communication system called Mod Mail. Subreddit leaders often prominently display their rules on the side of the subreddits’ main page and use built-in wiki features to document community guidelines and FAQ’s. Subreddit rules usually have to do with community topic or behavioral expectations (Fiesler et al., 2018). Studies in social computing have mapped out the kinds of rules that are adopted across subreddits (Chandrasekharan et al., 2018; Fiesler et al., 2018) and quantitative factors for rule change (Reddy and Chandrasekharan, 2023).
@@ 3.2. Fandom wikis
Wikis are collaboration-based websites where users peer produce public information repositories, like encyclopedias, by directly creating, editing, and organizing content on the website’s pages. Fandom is a platform that affords the creation and management of wikis, often made up of members of a fandom who share a strong interest in documenting information about a specific work of fiction or cultural artifact (e.g., TV show, book series, or video game). Interactions primarily occur on the “Talk Pages” of articles in a Fandom wiki, where members discuss article content. Many Fandom wikis activate a forum feature in the platform.
As the goal of wikis is to produce relevant, high-quality articles documenting their fandoms, leaders frequently develop rules regulating article content. Users who create a Fandom wiki become its de facto leader but frequently appoint other community members into leadership positions. Social computing scholars have documented how wikis engage in ongoing deliberation about how to change their communities’ rules (Keegan and Fiesler, 2017), and shown that as they grow and mature they become more bureaucratic (by adding more rules) (Butler et al., 2008) and oligarchic and closed to newcomers (Shaw and Hill, 2014; Halfaker et al., 2013; TeBlunthuis et al., 2018).
@@ 3.3. Fediverse servers
The Fediverse is an open, decentralized network of servers (also called “instances”) that run the ActivityPub protocol. Users join the Fediverse by registering an account on a server of their choice. The server enables its members to share and receive content from other servers, allowing users to interact with fellow server members and the broader network through comments, reposts (“boosts”), likes, and more. We focused on platforms using the Mastodon and Pleroma software, which both offer a microblogging experience similar to systems like Twitter. Servers we observed were generally organized around a shared interest or identity, but members were often not required to post content related to that theme.
Envisioning decentralized social media ecosystems without centralized owners, Fediverse servers using Mastodon and Pleroma are governed exclusively by the servers’ creators, who are referred to as “admins.” Admins control membership through open, closed, or application-based user registration, moderate content through admin panels and back-end access to data, and can block other servers to prevent interactions between their server and another (known as “defederation”). Rules on Fediverse servers are usually about speech, especially civility and legality. Although scholarly attention to Fediverse is relatively nascent, researchers have examined migrations to decentralized platforms (Dunbar-Hester, 2024; Cava et al., 2023; He et al., 2023; Andrade Nunes, 2023; Jeong et al., 2024), how platforms conceptualize decentralization (Zulli et al., 2020; Gehl and Zulli, 2023; Raman et al., 2019), community governance challenges (Anaobi et al., 2023; La Cava et al., 2024; Bono et al., 2024; Hwang et al., 2025; Zhang et al., 2024), coordination against hate speech (Caelin, 2022), rules on Mastodon platforms (Nicholson et al., 2023), and the effects of defederation on activity (Colglazier et al., 2024).
@@ 4.1. Community Sampling and Participant Recruitment
To investigate the origins of rules, we sought a sample of communities that varied in terms of goals. Between 2020 and 2021, we built datasets of community rules from subreddits, Fandom wikis, and Fediverse servers. Because we were initially motivated by observations of rule similarity across communities (Hwang and Shaw, 2022; Fiesler et al., 2018; Reddy and Chandrasekharan, 2023; Nicholson et al., 2023), we computationally clustered communities within each system based on textual similarities between their rules. We then screened clusters to identify a diverse, theoretically relevant sample of active communities to target for interviews. A more detailed description of this process can be found in Appendix Community Sampling Methods.
@@ 4.2. Interviews
We conducted semi-structured interviews via video calls between June and September 2022. All participants were either volunteer moderators or administrators familiar with their community’s rule history and above the age of 18. Interviews ranged from 30 to 120 minutes in length, averaging 60 minutes. A combination of trained undergraduate research assistants and the paper authors conducted the interviews following a protocol that asked participants to describe or tell us stories about their interactions with rules in their communities (see Appendix C). We asked questions about the origins of their rules, how they were enforced, when they changed and why, and about related challenges in governing online communities.
@@ 4.3. Analytical Process
Our analytical process was based on Charmaz’s approach to grounded theory (Charmaz, 2015), thematic analysis (Braun and Clarke, 2006), and collaborative affinity mapping. Small teams of trained research assistants led by four authors conducted line-by-line coding of interview transcriptions using the Taguette software, inductively coding text to break down participant’s experiences into actions and feelings through gerunds (e.g., “relaxing the rules” or “feeling a sense of belonging”). Together, we developed over 4,500 open initial codes. Discussions of these codes across the authorship team led to a rough list of 20 themes. Guided by this list, three authors then conducted a round of focused coding, a step which merges similar initial codes together and elevates them with broader conceptual labels. This led to 337 unique focused codes.
Three authors then collaboratively engaged in an in-depth affinity mapping exercise with the 337 focused codes to visualize patterns, themes, and processes via a digital whiteboard. Rounds of memo-writing and weekly meetings for discussion and iteration complemented this process. The three authors reached agreement on the most significant contributions to bring forward in this paper: rule adoption and rule change. Finally, the first author engaged in a round of theoretical sampling by using the list of 337 focused codes as a codebook to apply to all 40 interview transcripts. This was to identify supporting evidence across all three empirical settings and to analyze differences when patterns varied between them. The first author distilled a core set of processes and beliefs around rule adoption and change, which was iterated upon with feedback and input from the other authors.
@@ 5.1. Searching for Rule Templates
Few leaders we spoke to imagined a rule set during community creation. Instead, most leaders described the action of searching for rule templates in other online communities. For example, FA14, the leader of a toy franchise’s Fandom wiki, told us that they “are always inspired by other communities” and “always take a template of rules” from other communities. Some leaders described feelings of not being “actually experienced enough and knowledgeable enough” (FV12) to write rules themselves. FV10, a Fediverse leader for a regional community in the U.S., told us how “a big chunk” of their rules were made at community creation, saying that they “looked to other examples of rules” across Fediverse communities. They referenced their legal background to explain choosing templates over drafting rules themselves:
@@ 5.1.1. Regulating Content and Behavior
In the process for searching for rule templates, leaders often cited two motivations: regulating content and regulating behavior. For subreddit leaders like R1, rules regulating content were important for facilitating meaningful discussions. R1 explained why their community’s rules about mandating substantial discussion posts about a specific television genre were important:
@@
Leaders on subreddits often implemented rules on post formatting and content. For example, some communities had strict rules on advertising, such as R4’s IT professional discussion-based community. Others had regulations on content they called “misinformation,” such as R5’s regional community, which had rules adopted to prevent the spread of conspiracies related to the COVID-19 pandemic. We found that Fediverse servers rarely required specific kinds of content in their rules. However, they did restrict kinds of content that are inappropriate for professional work settings (NSFW content), hateful content, and advertising.
Fandom leaders view their rules as setting standards. For FA6, a leader of community for a multiplayer video game, rules governed information presentation and required content to be accurate. This was typical of rules across the Fandom wikis, almost all of which have content requirements and restrictions on formatting, article relevance, and accuracy.
Behavioral regulation was another important justification for rule adoption across all three settings. For subreddit leaders, this was expressed by a desire to ensure discussion spaces remained civil and safe. R9, a leader of a subreddit that organizes trades of goods for a specific hobby, summarized the sentiment that we heard from other subreddit leaders: “I think the biggest purpose of the rules is to keep everybody safe—to make sure nobody’s getting scammed…to make it so the community doesn’t become a toxic place.”
Although Fandom wikis place less emphasis on facilitating discussion, leaders still adopted rules to promote civility in Talk Pages, discussion forums, and in affiliated Discord servers that many Fandom communities used for synchronous chat. FA11 summarized the justification for behavioral rules we heard from other Fandom wiki leaders: “They’re there for safety, and they’re also there to make people feel like they’re included and they belong.”
@@ 5.2. Copying Other Communities’ Rules
One of the original questions motivating our study was “Why do so many online communities have similar rules?” In the previous section, we showed how leaders would seek out other communities’ rule sets to use as starting templates. For many, the next step was to copy. Copying was most evident in our interviews with members of the Fediverse. FV10 explained this step as “copying the good stuff” after they “look[ed] at other examples of how people have done it.” Leaders would copy and paste rules they felt were a good match for their community and then later “sort of build on it” (FV10) with further customization to their communities’ unique needs. That said, many leaders from the Fediverse would copy the entire rule set of another community and never customize or change it. In explaining these actions, these leaders emphasized a relational function of rules with the goal of establishing external legitimacy among other communities in their decentralized networks.
@@ 5.2.1. Establishing External Legitimacy
A key motivation behind the process of copying other communities’ rules was the desire to project legitimate governance (i.e., the presence of responsible content moderation) and communicate alignment with established norms across an ecosystem of online communities. Fediverse leaders have the power to block (“defederate”) communities that threaten their well-being and goals. These same leaders felt their choices were being observed and judged by admins of other Fediverse servers that might need to consider whether to defederate at some point. FV10 said this adds “an important function of rules” because it “sends signals to other admins that they know what we are moderating.” This external scrutiny from other admins motivated Fediverse leaders to adopt rules because these rules would signal accountable governance and content moderation. In this way, rules help establish external legitimacy for communities in decentralized networks.
Many Fediverse leaders pointed to mastodon.social’s rule set as a benchmark of legitimate community governance. mastodon.social was one of the first communities on the Fediverse and was created by Eugen Rochko, the developer of the most widely used platform for hosting Fediverse communities (Mastodon). Fediverse leaders often looked to this community because it’s “the largest instance” and because it’s “hosted by Eugen” (FV5). FV5 further explained, “You know, if you’re gonna copy rules about a new community, you might as well go right from the source. So I copied most or all of the rules.” The “source” in this case was mastodon.social. FV2 came to a similar conclusion:
@@
FV2 observed many other communities on the Fediverse drawing on mastodon.social’s rule set and this helped rationalize adopting it as well. We can infer from interview with FV2 that widespread adoption of mastodon.social’s rules happened in a self-reinforcing dynamic—communities with these rules signal their alignment with the standards and norms of a reputable, established community.
@@
Subreddit and Fandom wiki leaders did not mention copying other communities rules to signal legitimacy in the way that Fediverse leaders did. That said, some Fandom Wiki leaders like FA3 sought out rules from “other big wikis” to see how others addressed similar issues and used their rules as starting points. Subreddit leaders like R4 and R5 mentioned copying guidelines from the “Reddiquette,” which is an informal set of behavioral guidelines written by other Reddit leaders similar to the FFMC.
@@ 5.3. Customizing the Adopted Rule Set
Customizing the adopted rule set was a natural next step for many leaders. After all, the communities from which they found their initial rule sets often differ in their specific goals. Several interviewees from Fandom and Reddit offered insight into this process. FA1, a leader of a Fandom wiki documenting fashion trends, explained that “it would be nice” if every other Fandom wiki had the same rules, but that rules have to be modified to each community’s unique needs: “every community is different.” Leaders often searched for specific kinds of rules from different communities to solve their problems. FA4, the leader of a Fandom wiki for a comic book franchise, described this process in the context of creating a rule on image usage by customizing the image policy from another wiki:
@@
We note that changes were not driven purely based on differences in community topic. Leaders also described trying to improve the language or normative effect of a rule. For example, R4 based their rules on other communities they had led and made them “better to a certain degree.” Sometimes leaders described these edits to other rules as minimal and changing them “a little bit.” (FV12) Other times, the process of rewriting the rules could be “agonizing” (FV11) as leaders struggled to make their rules “well-written.”